# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ 7 Conclusion
In this manuscript, we introduced FedMerDel, a Bayesian mixture modelling algorithm using variational inference and ‘global’ merge moves to enable federated learning without sharing the full dataset across nodes. This model showed good performance on both simulated and real-world data, with only slight reductions in performance compared to analysing the full dataset on one machine (when this is possible), and is particularly useful when federated learning is required.
The distribution of data into separate batches for initial inference before extending the merge move to a ‘global merge’, directly addressed both the challenge of scaling model-based clustering, and the challenge associated with clustering data which is decentralised and held in separate locations. This is important in scenarios where data privacy is a concern, such as with patient health records. We demonstrated that our method was effective in investigating multimorbidity in one anonymised EHR dataset. We showed that FedMerDel improved time efficiency compared to stepwise parallelisation of the usual variational EM algorithm, especially as dataset sizes increased. We further showed that variational distributions can be used to generate observations from a certain cluster.
There is clearly scope for improvement in some aspects of the model. For example, to scale up to even larger datasets, we could add extra steps in the global merge as in the SIGN model (Ni et al., 2019), where local clusters are combined over multiple steps over multiple computing cores.
@@
While we presented our approach in the context of modelling large categorical datasets, we note that it can be straightforwardly adapted to mixtures of Gaussians and other members of the exponential family. The use of sufficient statistics to allow for efficient ELBO calculation in global merges could potentially extend to any exponential family distribution; (Hughes & Sudderth, 2013) provides some further details.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
@@ Appendix A Variable Selection
As explored in Law et al. (2004) and Tadesse et al. (2005), we can introduce binary feature selection indicators $\gamma_{j}$ feature saliencies - where $\gamma_{j}=1$ if and only if the $j$-th covariate is relevant to the clustering structure, and irrelevant variables have their feature saliencies reduced to zero. The probability density for a data point $x_{n}$ in cluster $k$ is given by:
$f({\bf x}_{n}|\Phi_{k})$ $=\prod_{j=1}^{P}f_{j}(x_{nj}|\Phi_{kj})^{\gamma_{j}}f_{j}(x_{nj}|\Phi_{0j})^{1-\gamma_{j}}$ (16)
$\Phi_{0j}=[\phi_{0j1},...,\phi_{0jL_{j}}]$ are parameter estimates obtained for covariate $j$ under the assumption that there exists no clustering structure in the $j$-th covariate. These are precomputed using maximum likelihood estimates as seen in Savage et al. (2013). The priors associated with $\gamma_{j}$ for $j=1,...,P$ are as follows:
@@
Appendix B Hyperparameters
We detail here further parameter settings for the priors detailed in Section 2.1 of the main paper.
  • $\boldsymbol{\phi}$: $\epsilon_{jl}=1/L_{j}$ for each $j=1,...,P,l=1,...,L_{j}$. There is no prior favouring for a variable to take a certain value in any of the clusters.
  • $\boldsymbol{\gamma},\boldsymbol{\delta}$: We set the hyperparameter for $\delta$ to be $a=2$ in all simulations, as in Rao & Kirk (2024). The use of a hyperprior for $\gamma$ allows the prior probability of the inclusion of each covariate to be a target for inference. In the variational algorithm, intially we include all variables by setting $c_{j}=\mathbb{E}_{\gamma}(\gamma_{j})=1$ for all $j=1,...,P$. Variables are found to be irrelevant as the algorithm runs.
Appendix C Mean-Field Variational Inference (VI)
To improve the readability and continuity of the supplementary material, some content from the main manuscript is repeated in this section. The true Bayesian posterior distribution is intractable as it is computationally intractable to evaluate the marginal likelihood $p(X)$. We employ a variational inference (VI) approach to seek an approximate distribution $q$ which is as close as possible to the true posterior. We do this by optimising the Evidence Lower Bound (ELBO) given below, $\mathcal{L}(q)$, with respect to $q(\theta)$, where $\theta$ is a collection of all parameters in the model.
[…]
This is equivalent to minimising the Kullback-Leibler (KL) divergence between $q(\theta)$ and the true posterior, $p(\theta|X)$. We constrain $q$ to be a mean-field approximation which can be fully factorised as (in the variable selection setting): $q(\theta)=q_{Z}(Z)q_{\pi}(\pi)q_{\Phi}(\Phi)q_{\gamma}(\gamma)q_{\delta}(\delta)$. The mean-field approximation is simple enough for tractable approximation of the ELBO, and is sufficiently flexible to provide good approximations.
In mean-field VI, the optimal solution $q_{j}^{\ast}(\theta_{j})$ for each component of $\theta$, $\theta_{j}$, satisfies (Bishop, 2006):
[…]
We adopt an iterative procedure to optimise $\mathcal{L}(q)$ by cycling between optimising $q(\theta)$ with respect to each parameter in turn. This numerical procedure is analogous to an Expectation-Minimisation (EM) algorithm, and the variational Bayes algorithm can be reduced to the EM algorithm by setting the variational densities to point estimates (Neal & Hinton, 1998; Ghahramani & Beal, 2000). As the ELBO is convex with respect to each factor $q_{j}^{\ast}(\theta_{j})$, convergence of the algorithm is guaranteed (Boyd & Vandenberghe, 2004). The ELBO is calculated after every EM step, and we can use sufficient statistics (similar to those detailed in Section 4.2) to reduce computation in sums over $N$ in the ELBO. This improves the efficiency of the overall algorithm. It is important to note that the frequency of ELBO calculations is a factor contributing to the overall run-time of any variational algorithm.
‘E’ Step
In the variational ‘E’ (expectation) step, the latent variables $Z$ are updated using the current variational distributions of all other parameters ($q^{\ast}(\pi)$, $q^{\ast}(\phi)$ and $q^{\ast}(\gamma)$, $q^{\ast}(\delta)$ in the variable selection case). Using Equation 20, in the model without variable selection, we have:
$r_{nk}$ is the responsibility of the $k$-th component for the $n$-th observation, and $\mathbb{E}[z_{nk}]=r_{nk}$. Data point $n$ is allocated to cluster $k^{\ast}$ where $k^{\ast}=\operatorname*{arg\,max}_{k}r_{nk}$.
‘M’ Step
Given updated responsibilities $r_{nk}$ in the ‘E’ step, the variational ‘M’ (maximisation) step updates cluster-specific parameters (and variable selection parameters) in the model.
where for $k=1,...,K,j=1,...,P,l=1,...,L_{j}$:
$\alpha^{\ast}_{k}$ $=\alpha_{0}+\sum_{n=1}^{N}r_{nk}$ (25)
$\epsilon^{\ast}_{kjl}$ $=\epsilon_{j}+\sum_{n=1}^{N}\mathbb{I}(x_{nj}=l)r_{nk}$ (26)
Expectations throughout are taken with respect to variational distributions. Between each ‘E’ and ‘M’ step, we calculate the value of the ELBO and use this to monitor convergence. All values are initialised using k-modes (Chaturvedi et al., 2001), a method analogous to k-means for categorical data.
Appendix D Variational Updates with Variable Selection
@@
All expectations throughout are taken over the variational distributions for each parameter. The general algorithm involves cycling between performing the ‘E’ step to evaluate the values of $r_{nk}$ with the current parameter values, and then the ‘M’ step to re-optimise all cluster-specific parameters and variable selection parameters.
Appendix E Merge and Delete Moves: Further Detail
E.1 Merge Moves
@@
We then perform a ‘dummy’ variational M step, updating the $\pi,\phi$ parameters. We run an extra E and an extra M step to allow observations to move in or out of the new cluster. We then calculate the ELBO and compare this to the ELBO before the merge move; we accept this move if the ELBO has improved, otherwise we return to the model prior to the merge. If accepted, we reassign cluster $k^{\ast}$ to the $k_{1}$ position and delete cluster $k_{2}$ from the model.
E.1.1 Selecting Candidate Clusters to Merge
We could choose a pair of clusters for a merge at random, but a merge is more likely to be accepted if the two clusters are ‘similar’. Making informed decisions on which pair of clusters to propose for a merge can improve efficiency by reducing the number of rejected merge moves. Two methods of assessing similarity between clusters using correlation and divergences between variational distributions are detailed below. Other ways include comparing marginal likelihoods or posterior probabilities of candidate models as in Hughes et al. (2015) and Ueda et al. (2000).
Due to the stochastic nature of the selection from the ‘candidate clusters’, similar results between different merge candidates are to be expected and the choice of measure is insignificant. Results from a simulation study comparing selection criteria for clusters to be merged are seen in Appendix J.4. Correlation and divergence-based methods performed similarly, but did provide a reduction in overall computation time and an improvement in rates of accepted merges compared to random selection.
@@ E.1.1 Selecting Candidate Clusters to Merge
We can use an assessment of correlation between the variational parameters associated with each cluster. For example, in Hughes & Sudderth (2013), two topics in the hierarchical Dirichlet process are proposed for a merge if the correlation between statistics representing usage of the topic across documents is above 0.05.
We can construct a correlation score by looking at the correlation between the parameter values $\epsilon^{\ast}$ in the distributions $q(\phi_{kj})$. In the binary case, each $\epsilon^{\ast}_{kj}$ is given by a two dimensional vector, and without loss of generality, for each k, we can take the first value of each of these vectors for $j=1,...,P$, concatenate these into one $j$ length vector, $\epsilon^{corr}_{k}$ and then compare the correlation of these concatenated vectors across clusters. The $\epsilon^{corr}_{k}$ values represent the (expected) frequency of 0’s in each cluster across all variables in a binary dataset of 0’s and 1’s.
We select randomly between the 3 cluster pairs with the highest correlation, provided $\mbox{Corr}(\epsilon^{corr}_{k_{1}},\epsilon^{corr}_{k_{2}})$ is above 0.05. With the categorical case with $L$ categories per variable, we could consider finding the correlation scores for $L-1$ concatenated vectors across all variables for $L-1$ of the categories (avoiding overparameterisation) and taking the mean correlation score for comparison. However, this is complicated when variables have different numbers of categories, and this method is less straightforward to extend to other probability distributions.
@@
We also look at using measures of distance between probability distributions to quantify similarity between clusters. For a pair of clusters $k_{1}$ and $k_{2}$, we look at the divergence between $q(\phi_{k_{1}j})$ and $q(\phi_{k_{2}j})$ and sum over all variables $j$ for a pair of clusters $k_{1}$ and $k_{2}$. We then choose randomly between the 3 cluster pairs with the lowest divergences, for example.
Different divergences could be considered; in this paper we considered both the Kullback-Leibler (KL) divergence and the Bhattacharyya distance. We emphasise that we are not seeking an exact, precise characterisation of the ‘distance’ but are only making approximate comparisons between candidate cluster pairs.
E.2 Delete Moves
@@
Delete moves refer to moves to delete unnecessary extra clusters in the model. Redundant small clusters often remain at local optima in variational algorithms for mixture models. These clusters are usually inconsistent across different initialisations, suggesting that moves to remove them and reassign observations to other clusters could improve performance.
@@
Let cluster $k$ be a candidate cluster for deletion. For the set of observations in cluster $k$, ie. $S_{k}:=\{n=1\ldots N:z_{nk}=1\}$, we create a new data-frame $X_{-k}$ with the rows from observations in $S_{k}$ removed, and the column for cluster $k$ removed. We first perform the variational E step after removing all variational parameters relevant to cluster $k$ in the model. This recalculates all responsibilities for the remaining observations. We then perform the variational M step, which updates the $\pi,\phi$ (and $\gamma,\delta$ if variable selection is used) parameters.
@@
We then return to the original dataframe $X$ and then run an E step using $X$ with the current model parameters, which reassigns all observations in $X$ to new clusters. After performing another M step, we then calculate the value of the ELBO function. We accept the delete move if the new ELBO is higher.
@@
Note that in both the merge and delete moves described above, there is an extremely small approximation involved in setting $r_{nk}=0$ exactly for responsibilities associated with clusters ‘removed’ in merge/delete moves. We go into further detail in Appendix H, showing that this approximation simply replaces numbers extremely close to 0 (numbers smaller than $10^{-80}$) during the algorithm and has virtually no effect on the clustering structure.
E.2.1 Selecting Candidate Clusters to Delete
@@
Smaller clusters are likely to be better candidates for deletion as it is likely these observations have been ‘left out’. Extremely small clusters are also less informative in practical applications and may be less reproducible in other similar independent datasets. In this paper, we propose a cluster chosen randomly between all clusters which are smaller than 5% of the dataset; if there are none, we choose randomly between the three smallest clusters. As in Appendix E.1.1, the stochasticity of this process means that other methods should not give significantly different results.
Appendix F Global ELBO Calculation - Further Details
@@
where all expectations are taken with respect to the variational distributions for $Z,\pi,\Phi$. From standard properties of the Dirichlet distribution, we initially calculate:
$\mathbb{E}_{\pi}[\ln{\pi}_{k}]$ $=\psi(\alpha^{\ast}_{k})-\psi(\sum_{k=1}^{K}\alpha^{\ast}_{k})$ (37)
$\mathbb{E}_{\Phi}[\ln\phi_{kjl}]$ $=\psi(\epsilon^{\ast}_{kjl})-\psi(\sum_{l=1}^{L_{j}}\epsilon^{\ast}_{kjl})$ (38)
where $\psi(x)$ refers to the digamma function.
As written in the main text of the paper, we also use the current $\alpha^{\ast}$ and $\epsilon^{\ast}$ values to calculate the values of sufficient statistics:
@@
Each term in Equation (F) is as follows:
$\mathbb{E}_{Z,\Phi}[\ln p(X|Z,\Phi)]$ $=\mathbb{E}_{Z,\Phi}[\sum_{n=1}^{N}\sum_{k=1}^{K}{z_{nk}}(\sum_{j=1}^{P}\ln\phi_{kjx_{nj}})]$ (41)
$=\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}(\sum_{j=1}^{P}\mathbb{E}_{\Phi}[\ln\phi_{kjx_{nj}}])$ (42)
$=\sum_{n=1}^{N}\sum_{k=1}^{K}\sum_{j=1}^{P}\sum_{l=1}^{L_{j}}r_{nk}\mathbb{I}(x_{nj}=l)\mathbb{E}_{\Phi}[\ln\phi_{kjl}]$ (43)
$=\sum_{k=1}^{K}\sum_{j=1}^{P}\sum_{l=1}^{L_{j}}S_{kjl}\mathbb{E}_{\Phi}[\ln\phi_{kjl}]$ (44)
$\mathbb{E}_{Z,\pi}[\ln p(Z|\pi)]$ $=\mathbb{E}_{Z,\pi}[\sum_{n=1}^{N}\sum_{k=1}^{K}{z_{nk}}\ln\pi_{k}]$ (45)
$=\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\mathbb{E}_{\pi}[\ln\pi_{k}]$ (46)
$=\sum_{k=1}^{K}T_{k}\mathbb{E}_{\pi}[\ln\pi_{k}]$ (47)
@@
Apart from the assignment entropy term in Equation (50), as we have calculated $T_{k}$ and $S_{kjl}$ via simple linear operations on our model parameters, there are no sums over $N$ and all terms are mainly sums over much smaller $K$, $P$, $L_{j}$. Furthermore, there are terms in $B(\alpha)$, $B(\alpha^{\ast})$, $B(\epsilon_{kj})$, $B(\epsilon^{\ast}_{kj})$ which cancel out - see Appendix H for more details. Importantly, $X$ appears nowhere in the calculations; our global merge move allows us to seek a global clustering structure without any full scan over the $N\times P$ matrix $X$.
For the assignment entropy term $\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\ln r_{nk}$, as detailed in the main paper, we can either calculate this value in parallel across different batches of observations across different cores, or in the case of the greedy search, this value remains constant as the merge moves only allow merges between clusters from different batches.
Appendix G Global Merge - Variable Selection

### v2

@@ 7 Discussion
In this manuscript, we introduced FedMerDel, a Bayesian mixture modelling algorithm using variational inference and ‘global’ merge moves to enable federated learning (FL) without sharing the full dataset across nodes. This model showed good performance on both simulated and real-world data, with only slight reductions in performance compared to centralised analysis, and excelled when data was not identically distributed across batches, a key challenge in FL. FedMerDel outperformed other existing unsupervised FL methods, where it could directly address both the challenge of scaling model-based clustering, and the challenge of clustering decentralised data. This is valuable for scenarios requiring both data privacy and scalability e.g. patient health records.
There is clearly scope for improvement in some aspects of the model. While we push scalability beyond other probabilistic approaches (eg. MCMC, EM), physical hardware limits remain, with memory use from large parameter sets creating computational bottlenecks. To scale further, we could add extra steps in the global merge as in the SIGN model (Ni2019), where local clusters are combined over multiple steps over multiple computing nodes or cores. Efficiency could be improved by adapting the ‘laps’ parameter controlling the frequency of merge/delete moves as MerDel approaches convergence, similar to learning rate scheduling. Future work could explore allowing local cluster splits during global merging (Song2020-qb represents local clusters as a mixture of distributions) and theoretical convergence guarantees for global merge moves. Further VI techniques including Stochastic Variational Inference (SVI) (JMLR:v14:hoffman13a) and variational tempering (pmlr-v51-mandt16) could also be implemented within each data batch to further improve scalability and accuracy of the local clustering.
@@
While we presented our approach in the context of modelling large categorical datasets, we note that it can be straightforwardly adapted to mixtures of Gaussians and other members of the exponential family for other applications. The use of sufficient statistics to allow for efficient ELBO calculation could extend to any exponential family distribution; hughes_memoized_2013 provides further details.
[…]
The project was part of the Bringing Innovative Research Methods to Clustering Analysis of Multimorbidity (BIRM-CAM) project that was funded by the MRC (MR/S027602/1). We also acknowledge MC_UU_00040/05.
@@ Appendix A Variable Selection
As explored in Law2004 and Tadesse2005, we can introduce binary feature selection indicators $\gamma_{j}$ - feature saliencies - where $\gamma_{j}=1$ if and only if the $j$-th covariate is relevant to the clustering structure, and irrelevant variables have their feature saliencies reduced to zero. The probability density for a data point $x_{n}$ in cluster $k$ is given by:
$f({\mathbf{x}}_{n}|\Phi_{k})$ $=\prod_{j=1}^{P}f_{j}(x_{nj}|\Phi_{kj})^{\gamma_{j}}f_{j}(x_{nj}|\Phi_{0j})^{1-\gamma_{j}}$ (14)
$\Phi_{0j}=[\phi_{0j1},...,\phi_{0jL_{j}}]$ are parameter estimates obtained for covariate $j$ under the assumption that there exists no clustering structure in the $j$-th covariate. These are precomputed using maximum likelihood estimates as seen in Savage2013. The priors associated with $\gamma_{j}$ for j=1,,Pj=1,...,P are as follows:
@@
These irrelevant variables can correspond to noisy/corrupted features, which is particularly useful in real-world cases where we believe there may be data corruption.
Appendix B Variational Updates with Variable Selection
@@
As before, all expectations throughout are taken over the variational distributions for each parameter and we alternate between ‘E’ and ‘M’ steps.
Appendix C Merge and Delete Moves: Further Detail
C.1 Merge Moves
@@
We then perform a ‘dummy’ variational M step, updating the $\pi,\phi$ parameters. We run an extra E and an extra M step to allow observations to move in or out of the new cluster. We calculate the ELBO and compare this to the ELBO before the merge move; we accept this move if the ELBO has improved, otherwise we return to the model prior to the merge. If accepted, we reassign cluster $k^{\ast}$ to the $k_{1}$ position and delete cluster $k_{2}$ from the model.
C.1.1 Selecting Candidate Clusters to Merge
We could choose a pair of clusters for a merge at random, but a merge is more likely to be accepted if the two clusters are ‘similar’. Making informed decisions on which pair of clusters to propose for a merge can improve efficiency by reducing the number of rejected merge moves. Two methods of assessing similarity between clusters using correlation and divergences between variational distributions are detailed below. Other ways include comparing marginal likelihoods or posterior probabilities of candidate models as in pmlr-v38-hughes15 and ueda_smem_2000.
Due to the stochastic nature of the selection from the ‘candidate clusters’, similar results between different merge candidates are to be expected and the choice of measure is insignificant. Results from a simulation study comparing selection criteria for clusters to be merged are seen in Appendix H.4. Correlation and divergence-based methods performed similarly, but did provide a reduction in overall computation time and an improvement in rates of accepted merges compared to random selection.
@@ C.1.1 Selecting Candidate Clusters to Merge
We can use an assessment of correlation between the variational parameters associated with clusters; a notion of correlation is seen also in hughes_memoized_2013.
We construct a correlation score by looking at the correlation between the parameter values $\epsilon^{\ast}$ in the distributions $q(\phi_{kj})$. In the binary case, each $\epsilon^{\ast}_{kj}$ is given by a two dimensional vector. For each k, we can take the first value of each of these vectors for j=1,,Pj=1,...,P, concatenate these into one $j$ length vector, $\epsilon^{corr}_{k}$ and then compare the correlation of these concatenated vectors across clusters. The $\epsilon^{corr}_{k}$ values represent the (expected) frequency of 0’s in each cluster across all variables in a binary dataset of 0’s and 1’s.
We select randomly between the 3 cluster pairs with the highest correlation, provided $\mbox{Corr}(\epsilon^{corr}_{k_{1}},\epsilon^{corr}_{k_{2}})$ is above 0.05. With the categorical case with $L$ categories per variable, we could consider finding the correlation scores for L1L-1 concatenated vectors across all variables for L1L-1 of the categories and taking the mean correlation score. However, this is complicated when variables have different numbers of categories.
@@
We also look at using measures of distance between probability distributions to quantify cluster similarity. For a pair of clusters $k_{1}$ and $k_{2}$, we look at the divergence between $q(\phi_{k_{1}j})$ and $q(\phi_{k_{2}j})$ and sum over all variables $j$ for a pair of clusters $k_{1}$ and $k_{2}$. We then eg. choose randomly between the 3 cluster pairs with the lowest divergences. Many divergences could be considered; in this paper we considered the Kullback-Leibler (KL) divergence and the Bhattacharyya distance. We emphasise that we are not seeking a precise characterisation of the ‘distance’ but are only making approximate comparisons.
C.2 Delete Moves
@@
Delete moves refer to moves to delete unnecessary extra clusters in the model. Redundant small clusters often remain at local optima in variational algorithms for mixture models.
@@
Let cluster $k$ be a candidate cluster for deletion. For the set of observations in cluster $k$, ie. $S_{k}:=\{n=1\ldots N:z_{nk}=1\}$, we create a new data-frame $X_{-k}$ with the rows from observations in $S_{k}$ removed, and the column for cluster $k$ removed. We first perform the variational E step after removing all variational parameters relevant to cluster $k$ in the model. This recalculates all responsibilities for the remaining observations. We then perform the variational M step, which updates cluster-specific parameters.
@@
We then return to the original dataframe $X$ and then run an E step using $X$ with the current model parameters, which reassigns all observations in $X$ to new clusters. After another M step, we then calculate the value of the ELBO function. We accept the delete move if the new ELBO is higher.
@@
Note that in both the merge and delete moves described above, there is a small approximation involved in setting $r_{nk}=0$ exactly for responsibilities associated with clusters ‘removed’. Appendix E gives further detail; this approximation simply replaces infinitesimally small numbers (smaller than $10^{-80}$) with 0.
C.2.1 Selecting Candidate Clusters to Delete
@@
Smaller clusters are likely to be better candidates for deletion as it is likely these observations have been ‘left out’. Extremely small clusters are also less informative in practical applications and may be less reproducible in other similar independent datasets. In this paper, we propose a cluster chosen randomly between all clusters which are smaller than 5% of the dataset; if there are none, we choose randomly between the three smallest clusters.
Appendix D Global Merge - Further Details
D.1 Random and Greedy Search
Greedy Search
We prevent merges between two clusters from the same batch with a greedy search, on the hypothesis that such merges would already have taken place at the local stage. This can be carried out in a sequential manner by examining the correlation/divergences between Cluster 1 of Batch 1 and all clusters in Batch 2. Once Cluster 1 has merged with any cluster in Batch 2, we move to Batch 3 and consider merging with the clusters in this batch. We repeat for all batches, before moving to Cluster 2 of Batch 1 and so on until we have gone through all clusters in Batch 1. We then repeat the process with Batches 2, 3, …, B, looking only at Batches b+1,,Bb+1,...,B for Batch $b$.
Random Search
Similarly to local merge moves, we calculate correlations/divergences between all $K$ clusters, and pick the merge candidate pair randomly between the 3 pairs with the highest correlation (provided correlation >> 0.05) or lowest divergence. There is no restriction on proposing merges between two clusters of the same batch. This stops after a fixed number of global merge moves (e.g. 10) have been rejected consecutively.
D.2 Global ELBO Calculation
@@
where all expectations are taken with respect to the variational distributions for $Z,\pi,\Phi$. We use the current $\alpha^{\ast}$ and $\epsilon^{\ast}$ values to calculate the values of sufficient statistics:
@@
Each term in Equation (D.2) is as follows:
$\mathbb{E}_{Z,\Phi}[\ln p(X|Z,\Phi)]$ $=\mathbb{E}_{Z,\Phi}[\sum_{n=1}^{N}\sum_{k=1}^{K}{z_{nk}}(\sum_{j=1}^{P}\ln\phi_{kjx_{nj}})]=\sum_{k=1}^{K}\sum_{j=1}^{P}\sum_{l=1}^{L_{j}}S_{kjl}\mathbb{E}_{\Phi}[\ln\phi_{kjl}]$ (29)
$\mathbb{E}_{Z,\pi}[\ln p(Z|\pi)]$ $=\mathbb{E}_{Z,\pi}[\sum_{n=1}^{N}\sum_{k=1}^{K}{z_{nk}}\ln\pi_{k}]=\sum_{k=1}^{K}T_{k}\mathbb{E}_{\pi}[\ln\pi_{k}]$ (30)
@@
Apart from the assignment entropy term in Equation (33), as we have calculated $T_{k}$ and $S_{kjl}$ via simple linear operations on our model parameters, there are no sums over $N$ and all terms are mainly sums over much smaller $K$, $P$, $L_{j}$. Furthermore, some terms in $B(\alpha)$, $B(\alpha^{\ast})$, $B(\epsilon_{kj})$, $B(\epsilon^{\ast}_{kj})$ which cancel out. Importantly, $X$ appears nowhere in the calculations; our global merge move allows us to seek a global clustering structure without any full scan over the $N\times P$ matrix $X$.
D.3 Global Merge - Variable Selection